what mass of CH4 formed when 8.3 moles of CO react with        hydrogen according to following equation?

2CO  +  2H2  → CH4 + CO2



1
Expert's answer
2021-06-18T07:21:29-0400

Solution:

Balanced chemical equation:

2CO + 2H2 → CH4 + CO2

According to the equation above: n(CO)/2 = n(CH4)

n(CH4) = n(CO) / 2 = 8.3 mol / 2 = 4.15 mol


The molar mass of CH4 is 16.04 g/mol.

Therefore,

(4.15 mol CH4) × (16.04 g CH4 / 1 mol CH4) = 66.566 g CH4 = 66.6 g CH4


Answer: 66.6 grams of CH4 formed.